Visit the Yahoo web site and click the link to Finance. Enter the letters "XOM" into the Quote Lookup box, and this will display the latest stock price, related stock data, and links to the financial statements for Exxon-Mobil Corporation ("XOM" is the stock symbol for this company). Using the available data, indicate the EPS and P/E ratios. Next, digging deeper into the financials, perform a horizontal analysis spanning the latest three years of data, specifically showing the trends (change over time) for earnings, inventory, the company's cash position, and also indicate the trend for the gross profit ratio and profit margin ratio. In addition, perform a vertical analysis over all of the asset classes listed on the Balance Sheet (as a proportion of the Total Assets). Summarize all of your findings, and be sure to display all calculations you've made in the process. Also provide conclusions regarding your assessment of what your analysis indicates about the current state of Exxon-Mobil.
Answer #1

1) Basic EPC is 4.63 and Diluted EPS is also 4.63

2)P/E Ratio = MPS/EPS

MPS of XOM-Exxon-Mobil Corporation on 31st March 2018 is Rs. 83.64/-

=83.64/4.63 =18.06 times2017 2016 2015 9.0 3.97.9 11.1 4.6 9.4 17.9 19.7 18.0 16.8 18.4 16.5 0.82 0.87 0.79 13.2 5.7 17.6 KEY FINANCIAL RATIOS Return on average capital employed (percent) Earnings to average ExxonMobil share of equity (percent) Debt to capital (percent) Net debt to capital (percent) Current assets to current liabilities (times) Fixed-charge coverage (times)